OpenStudy (anonymous):

The area of a square is 222 meters. What number is closest to the length of each side of the square?

OpenStudy (kropot72):

Area is not measured in meters. The square meter is a measure of area.

OpenStudy (anonymous):

Thats what I meant, sorry

OpenStudy (anonymous):

by closest number, do you mean closest whole number?

OpenStudy (anonymous):

closest fraction?

OpenStudy (anonymous):

The answer choices are A. 14.5 meters B. 14.9 meters C. 15 meters D. 16 meters

OpenStudy (anonymous):

those are pretty close together... especially b and c because it's one of em..: 14*14 = 196, nowhere close to 222 15*15 = 225, close to 222 14.9 * 14.9 =222.01 , ooh even closer!!!

OpenStudy (anonymous):

Lol, ok, I see. thanks

OpenStudy (anonymous):

yw...)